OpenStudy (anonymous):

r = 1 is a root of 4x^2 – 3x – 1 = 0 ^ this false? r = -1 is a root of 4x^2 – 3x – 1 = 0 ^ it is true?

OpenStudy (lgbasallote):

substitute -1 into x of 4x^2 - 3x - 1 \[\implies 4(-1) ^2 - 3(-1) - 1 \implies 4(1) + 3 - 1 \implies 4 + 3 - 1 \implies 6\] therefore r = -1 is NOT a root if it's a root..the result should be 0

OpenStudy (lgbasallote):

does that help?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

the first came as true for me?

OpenStudy (lgbasallote):

you mean r = 1?

OpenStudy (anonymous):

yeah

OpenStudy (lgbasallote):

yes it's a root becuase 4(1)^2 - 3(1) - 1 = 4 - 3 - 1 = 4 - 4 = 0